SPONTANEOUS EMISSION BETWEEN ORTHO- AND PARA-LEVELS OF WATER-ION, H_{2}O^{+}

\underline{KEIICHI TANAKA}, Department of Applied Chemistry, National Chiao Tung University, Hsinchu, 30010, TAIWAN, Department of Chemistry, Faculty of Sciences, Kyushu University, Fukuoka, 812-8581 JAPAN; KENSUKE HARADA, Department of Chemistry, Faculty of Sciences, Kyushu University, Fukuoka, 812-8581 JAPAN; SHINKOH NANBU, Department of Materials and Life Sciences, Faculty of Science and Engineering, Sophia University, Tokyo 102-8554, JAPAN; TAKESHI OKA, Department of Astronomy and Astrophysics and Department of Chemistry, the Enrico Fermi Institute, the University of Chicago, Chicago, Illinois, 60637, USA.

Nuclear spin conversion interaction of water ion, H_{2}O^{+}, has been studied to derive spontaneous emission lifetime between ortho- and para-levels. H_{2}O^{+} is a radical ion with the ^{2}B_{1} electronic ground state. Its off-diagonal electron spin-nuclear spin interaction term, T_{ab}(S_{a}ΔI_{b} + S_{b}ΔI_{a}), connects para and ortho levels, because ΔI = I_{1} − I_{2} has nonvanishing matrix elements between I = 0 and 1. The mixing by this term with T_{ab} = 72 MHz predicted by ab initio theory in the MRD-CI/Bk level,^{a} is many orders of magnitude larger than for closed shell molecules because of the large magnetic interaction due to the un-paired electron. Using the molecular constants reported by Mürtz et al. by FIR-LMR^{b}, we searched for ortho and para coupling channels below 1000 cm^{−1} with accidental near degeneracy between para and ortho levels. For example, hyperfine components of the 4_{2,2}(ortho) and 3_{3,0}(para) levels mix by 1.2 × 10^{−3} due to their near degeneracy (ΔE = 0.417 cm^{−1}), and give the ortho-para spontaneous emission lifetime of about 0.63 year. The most significant low lying 1_{0,1}(para) and 1_{1,1}(ortho) levels, on the contrary, mix only by 8.7 × 10^{−5} because of their large separation (ΔE = 16.267 cm^{−1}) and give the spontaneous emission lifetime from 1_{0,1}(para) to 0_{0,0}(ortho) of about 100 year.These results qualitatively help to understand the observed high ortho- to para- H_{2}O^{+} ratio of 4.8 ± 0.5^{c} toward Sgr B2 but they are too slow to compete with the conversion by collision unless the number density of the region is very low (n ∼ 1 cm^{−3}) or radiative temperature is very high (T_{r} > 100 K).

aM. Staikova, B. Engels, M. Perić, and S.D. Peyerimhoff, Mol. Phys. 80, 1485 (1993)

bP. Mürtz, L.R. Zink, K.M. Evenson, and J.M. Brown J. Chem. Phys. 109, 9744 (1998).

cLP. Schilke, et al., A&A 521, L11 (2010).